# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

### v3

@@ 4.1 Algorithm and Pseudocode
The core of our method is a four-step process that uses a sliding window to compare the geometric shape of a time series against itself.
@@
Second, for each window, we compute its corresponding path signature. The time-series data within the window is first normalised to ensure that the signature calculation is not sensitive to the absolute scale of the time or metric axes. Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$. The normalised two-dimensional path is then passed to the library, which computes the path signature up to a pre-specified truncation depth $d$. This operation yields two signature vectors, $S(W_{1})$ and $S(W_{2})$, which serve as high-dimensional feature representations of the performance trajectory within each window.
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors. This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows. A large distance implies a significant change in the underlying dynamics of the time series. To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance. In our retrospective evaluation we compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series for a given creative. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector. For online deployment, the same thresholding rule can be implemented using an initial calibration period or a rolling robust estimate of location and scale computed strictly from past distances.
Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments. For each segment, we perform a final trend classification. We fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope. This final step provides an interpretable label for the performance behaviour within each period of stability, making the results directly actionable for a marketing analyst. The complete process is summarised in Algorithm 1.
Algorithm 1 Signature-Based Change Point Detection
@@
4:for $i\leftarrow 1$ to $T-2w_{d}+1$ do
5:  $W_{1}\leftarrow Y[i:i+w_{d}]$
6:  $W_{2}\leftarrow Y[i+w_{d}:i+2w_{d}]$
7:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1], min-max scale metric to [0,1]
8:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
9:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
10:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
11:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
12:  Append $(Y[i+w_{d}].\text{date},D_{i})$ to $Distances$.
13:end for
@@
20:for each SegmentSegment in SegmentsSegments do
21:  $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
22:  Append TrendTrend to $T_{s}$.
23:end for
24:return $CP,T_{s}$
@@ 4.2 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w_{d}$ and the sensitivity threshold multiplier $k$. The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
The window size $w_{d}$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w_{d}$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
@@ 4.3 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w_{d}$ be the window size. The loop runs approximately $T-2w_{d}$ times. Inside the loop, the most expensive operation is the signature computation. For fixed path dimension $m=2$ and small truncation depth $d\in\{3,4\}$, signature computation is effectively linear in the number of time points per window. Formally, the number of truncated signature terms scales as $\sum_{k=0}^{d}m^{k}$, so computational cost grows quickly with $d$ but remains practical for small $d$. Since $w_{d}$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$. This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
5 Quantifying Wastage: The Financial Impact of Fatigue
A key advantage of our methodology is its ability to translate a statistical observation into a tangible financial metric. While identifying a change point is analytically valuable, its true business utility is realised when we can quantify the cost of inaction. To this end, we introduce a framework for calculating ’wastage’, defined as the opportunity cost incurred by continuing to invest in a creative asset after its performance has begun to degrade.
The calculation first requires the establishment of a fair performance benchmark. This benchmark represents the level of performance the creative was capable of achieving during its peak. We establish this by analysing the segments identified by the change point detection algorithm. The benchmark CTR, denoted $CTR_{bench}$, is calculated as the average CTR over the most effective period of stable or improving performance. This ensures the creative is judged against its own demonstrated potential rather than an arbitrary external standard.
With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance. To reduce sensitivity to day-to-day sampling noise, we compute lost clicks using the same noise-robust CTR estimate used in the onset definition below. This is given by the formula
@@ 5 Quantifying Wastage: The Financial Impact of Fatigue
where $\widehat{p}_{t}$ is the Beta–Binomial shrinkage estimate of CTR and $\text{Impressions}_{t}$ is the number of impressions on day $t$.
To translate these lost clicks into a monetary value, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period. The daily wastage is then
@@
The total wastage for a creative is the sum of this daily wastage over the entire period of underperformance. This framework provides a clear and interpretable financial metric that directly links the statistical detection of fatigue to a business-relevant outcome. It allows marketing managers to not only identify underperforming assets but also to prioritise them based on the financial impact of their decline, leading to more efficient and data-driven budget allocation.
@@ 6.1 Synthetic Data Framework
To address the confidentiality constraints inherent in digital advertising data while enabling methodological exposition and evaluation, we use a synthetic panel dataset designed to mimic common characteristics of performance marketing time series: (i) high-frequency daily measurement, (ii) substantial day-to-day volatility in CTR, and (iii) heterogeneous trajectories across creatives.
@@ 6.1.1 Dataset structure and provenance
The synthetic dataset is included with the repository (data/synthetic/synthetic_data.csv). It contains 3,077 observations from 2025–01–01 to 2025–06–30 across 17 creatives spanning multiple campaigns. Each record contains impressions and clicks (allowing CTR to be treated as a noisy binomial rate), alongside auxiliary spend fields. Most creatives have near-daily coverage over the study window, but a subset are intentionally irregularly observed to reflect real-world missingness (e.g., pacing, pauses, or delivery gaps). The CSV also includes a pre-computed CTR column; throughout this paper we define $CTR_{t}:=C_{t}/I_{t}$ and use clicks and impressions for all modelling and evaluation.
@@ 6.1.2 Alignment with empirical magnitudes
The panel is calibrated to resemble plausible ranges reported in industry and marketing analytics contexts: impression volumes are typically in the thousands to tens of thousands per day, and CTR values concentrate in the low single-percent range (approximately 0.5%–3.5%), consistent with broad digital advertising benchmarks (IAB, 2023). The dataset is not intended to reproduce any specific platform or sector; rather, it provides a transparent setting for comparing detectors under a clearly specified operational ground truth.
Importantly, the dataset does not include injected changepoints or labelled fatigue onsets. We therefore define fatigue onset operationally in Section 6.4 and evaluate methods against that rule-based definition.
@@ 6.2 Comparative Method Analysis
We benchmark our signature-based approach against established methods from both marketing practice and statistical literature. Table 6.2 provides a comprehensive comparison:
@@
Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No Tunable High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(n\,w_{reg})$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
@@
The signature-based approach offers several key advantages. Unlike machine learning methods, it requires no training and operates immediately on new creatives. Its responsiveness is tunable through the window size and threshold parameter, enabling a practical trade-off between early warning and alert burden. Its linear complexity enables scaling to portfolios containing thousands of creatives. Furthermore, the geometric robustness of the approach allows it to capture complex patterns beyond simple mean or variance shifts, making it particularly suited to the dynamic nature of advertising performance data.
@@ 6.3 Experimental Design
    Operational Ground Truth: Define fatigue onset $\tau$ per creative via the noise-robust, benchmark-relative rule in Section 6.4.
@@
    Method Application: Apply the signature-based detector and two baseline heuristics (moving-average crossover; rolling regression) to each creative.
@@
    Sensitivity Analysis: Vary detector parameters to characterise the trade-off between lead time and alert burden.
For the baseline methods, we use $CTR_{t}:=C_{t}/I_{t}$ and define windows over observed days (i.e., the last $k$ available observations for a creative) to accommodate missingness. Moving-average crossover uses trailing simple moving averages $MA_{7}(t)$ and $MA_{21}(t)$ computed over the last 7 and 21 observed days, respectively. An alert triggers on observed day $t$ if $MA_{7}(t)<MA_{21}(t)$ and this inequality holds for three consecutive observed days. Rolling regression fits an OLS model of $CTR$ on time over the last $w_{reg}=21$ observed days and triggers when the estimated slope is negative and statistically significant at $\alpha=0.05$ for three consecutive observed days.
The experimental framework ensures fair comparison by: - Using identical data for all methods - Using consistent pre-processing and smoothing assumptions in the operational ground truth - Defining a common alert date $\hat{\tau}$ for each method as specified in Section 7 - Reporting sensitivity analysis over detector parameters to characterise the early-warning versus alert-burden trade-off
6.4 Operational Definition of Fatigue Onset (Ground Truth)
Our empirical evaluation requires a reproducible definition of “fatigue onset” that is robust to the substantial day-to-day volatility inherent in digital advertising metrics. Because the dataset does not contain injected changepoints or labelled onset dates, we adopt an operational ground truth based on a sustained, economically meaningful deterioration in click-through rate (CTR) relative to a recent performance baseline.

### v4

@@ 4.3 Algorithm and Pseudocode
Algorithm 1 is the fatigue-specific online screen. The shape-only detector used in the panel experiment follows a related four-step process that uses a sliding window to compare the geometric shape of a time series against itself. It is included as an interpretable baseline and ablation rather than as the recommended fatigue-specific production rule.
@@
Second, for each window, we compute its corresponding path signature. The time-series data within the window is first normalised to reduce sensitivity to the absolute scale of the time or metric axes. Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$; flat metric windows are assigned a zero metric coordinate throughout. The normalised two-dimensional path is then passed to the library, which computes the path signature up to a pre-specified truncation depth $d$. This operation yields two signature vectors, $S(W_{1})$ and $S(W_{2})$, which serve as high-dimensional feature representations of the performance trajectory within each window. Because metric values are normalised within each window, this detector is intentionally invariant to absolute level and amplitude shifts; detecting those shifts requires a global scaling scheme or additional unnormalised features.
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D_{i}=\|S(W_{1}^{\text{norm}})-S(W_{2}^{\text{norm}})\|_{2}$, between the two signature vectors. This distance provides a scalar measure of dissimilarity in the transformed shape of the performance paths in the two adjacent windows. A large distance is treated as evidence of a candidate change in the underlying dynamics of the time series. To distinguish unusually large changes from background variation, we establish a threshold for this distance. In our retrospective panel evaluation we compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series for a given creative. This is an offline segmentation rule because early markers can depend on later distances. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D_{i}$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector. This is an anomaly threshold, not a formal significance threshold. For online deployment, the same thresholding idea must be implemented using an initial calibration period or a rolling estimate of location and scale computed strictly from past distances.
Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments. For each segment, we perform a final trend classification. Segments with fewer than three observations are labelled “stable/insufficient data” because the usual slope significance test has no positive residual degrees of freedom, even though a two-point slope can be computed. For segments with at least three observations, we fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope. This final step provides an interpretable label for the performance behaviour within each period of stability, making the results directly actionable for a marketing analyst. The complete process is summarised in Algorithm 2.
Algorithm 2 Shape-Only Signature Change Point Detection
@@
4:if $T<3$ then
5:  return $[]$, “stable/insufficient data”
6:end if
7:if $T<2w_{d}$ then
8:  return $[]$, $\text{ClassifyTrendViaRegression}(Y)$
9:end if
10:for $i\leftarrow 1$ to $T-2w_{d}+1$ do
11:  $W_{1}\leftarrow Y[i:i+w_{d}-1]$
12:  $W_{2}\leftarrow Y[i+w_{d}:i+2w_{d}-1]$
13:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1]; min-max metric or use zero for flat windows
14:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
15:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
16:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
17:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
18:  Append $(Y[i+w_{d}].\text{date},D_{i})$ to $Distances$.
19:end for
@@
26:for each SegmentSegment in SegmentsSegments do
27:  if $\text{Length}(Segment)<3$ then
28:   $Trend\leftarrow\text{``stable/insufficient data''}$
29:  else
30:   $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
31:  end if
32:  Append TrendTrend to $T_{s}$.
33:end for
34:return $CP,T_{s}$
@@ 4.4 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by the window size $w_{d}$, signature depth $d$, threshold multiplier $k$, and, for the level-aware fatigue screen, the directional-drop gate $\gamma$. In the injected benchmark we use $d=3$, $w_{d}=14$, $k=1.5$ for the headline level-aware online row, and $\gamma=2$.
The window size $w_{d}$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w_{d}$ depends on the cadence and volatility of the time series being analysed. The experiments below use daily windows in the 7–14 day range as exploratory operating points, not as universal defaults.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives through the rule
$D_{i}>\mu_{D}+k\sigma_{D},$
where $D_{i}$ is the adjacent-window signature distance and $\mu_{D},\sigma_{D}$ are computed from the relevant reference distribution of distances. In the retrospective panel experiment, the reference distribution is the full set of distances for that creative. In an online detector, it must be replaced by a past-only calibration distribution. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. In the experiments below, values between 1.5 and 2.5 are treated as exploratory operating points rather than universally recommended settings. The optimal choice should be tuned using validation data, explicit alert budgets, or historical data with known change points when available.
For synthetic-only deployment research, we recommend a fixed calibration protocol: choose $(w_{d},d,k,\gamma)$ on a synthetic training suite before evaluating new scenarios. The objective should be stated before fitting, for example maximising
$\text{mean captured proxy clicks per series}-c_{\mathrm{review}}\times\text{alert rate},$
where alert rate is the fraction of series with at least one alert under the evaluated rule. Alternatively, one can target a no-event false-alert rate under nuisance regimes. In Section 7.2 we implement the simplest version of this protocol by selecting the threshold setting on the first 12 replicates per scenario and evaluating the frozen setting on the remaining 12 replicates. This does not replace external validation, but it prevents the headline threshold from being justified only by the same rows used for evaluation.
@@ 4.5 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w_{d}$ be the window size. The loop runs approximately $T-2w_{d}$ times. Inside the loop, the most expensive operation is the signature computation. Formally, the full truncated signature has $\sum_{k=0}^{d}m^{k}$ terms, equivalently $\sum_{k=1}^{d}m^{k}$ non-constant terms if the level-0 term is omitted by the implementation. For $N$ admissible window positions, the signature-computation term is approximately $\mathcal{O}(Nw_{d}\sum_{k=1}^{d}m^{k})$. Thus, the detector scales linearly in $T$ only when the window size, path dimension, and truncation depth are treated as fixed constants. This fixed-depth, fixed-window regime is the one used in our experiments and is suitable for portfolio screening; production deployment still requires end-to-end load testing.
5 Quantifying Wastage: A Proxy for the Impact of Fatigue
One practical reason to pair detection with valuation is that a statistical alert is useful only if it helps prioritise action. To this end, we introduce a benchmark-CPC proxy for ’wastage’, defined as the opportunity value of clicks not obtained when a creative performs below its own benchmark. This is not an identified causal estimate of financial loss; it is a consistent valuation rule for ranking and triage.
The calculation first requires the establishment of a performance benchmark. For operational use, this benchmark must be computed using only data available up to the alert date, for example the best prior stable or improving segment identified before day $t$ or a detector-agnostic trailing recent-best period. The benchmark CTR, denoted $CTR_{bench}$, is defined on a comparable CTR scale using the same Beta-prior family as the daily estimator. One suitable definition is a posterior pooled benchmark over a benchmark period $B^{\star}$:
$CTR_{bench}=\frac{\sum_{u\in B^{\star}}C_{u}+\alpha}{\sum_{u\in B^{\star}}I_{u}+\alpha+\beta}.$
This impression-weighted construction aligns the benchmark with the subsequent lost-click calculation. Zero-impression days have no effect on the benchmark because they leave the pooled click and impression totals unchanged; the fixed prior contribution $\alpha+\beta$ is unaffected. The benchmark represents the level of performance the creative had recently demonstrated, not an externally imposed standard.
With the benchmark established, we can calculate the opportunity proxy for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate a smoothed ’lost-click’ proxy. This is not the literal realised-click shortfall $CTR_{bench}I_{t}-C_{t}$; it is a noise-stabilised proxy used for ranking and comparison. To reduce sensitivity to day-to-day sampling noise, we compute it using the same noise-robust CTR estimate used in the onset definition below. Specifically, for daily clicks $C_{t}$ and impressions $I_{t}$ we use the posterior mean
$\widehat{p}_{t}=\frac{C_{t}+\alpha}{I_{t}+\alpha+\beta},$
where $\alpha=m_{0}\bar{p}$, $\beta=m_{0}(1-\bar{p})$, $m_{0}=1000$, and $\bar{p}$ is the pooled CTR across the synthetic panel. In an operational online workflow, $\bar{p}$ must be fixed from historical training data available before the scored period; in this paper it is an offline synthetic-panel construction. Smoothed lost clicks are then
@@ 5 Quantifying Wastage: A Proxy for the Impact of Fatigue
where $\text{Impressions}_{t}$ is the number of impressions on day $t$.
To translate these lost clicks into a monetary proxy, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period. The daily proxy is then
@@
The total wastage proxy for a creative is the sum of this daily value over dates after the benchmark period for which $\widehat{p}_{t}<CTR_{bench}$. Alternative valuation rules could assign a different monetary value to lost clicks; the purpose here is to use a transparent and reproducible rule for comparing creatives and alert timings.
@@ 6.1 Synthetic Data Framework
To address the confidentiality constraints inherent in digital advertising data while enabling methodological exposition and evaluation, we use synthetic datasets only. This is a deliberate evidence boundary: the paper does not report client campaign histories, and it therefore does not claim external validation on production portfolios. The synthetic panel dataset is designed to mimic common characteristics of performance marketing time series: (i) high-frequency daily measurement, (ii) substantial day-to-day volatility in CTR, and (iii) heterogeneous trajectories across creatives.
@@ 6.1.1 Dataset structure and provenance
The synthetic dataset is included with the repository (data/synthetic/synthetic_data.csv). It contains 3,077 observations from 2025–01–01 to 2025–06–30 across 17 creatives spanning multiple campaigns, corresponding to a complete daily panel over this date range. Each record contains impressions and clicks, so for creative $i$ on observed day $t$ clicks can be modeled as $C_{it}\mid I_{it}\sim\mathrm{Binomial}(I_{it},p_{it})$, and the observed CTR $C_{it}/I_{it}$ is a noisy binomial proportion with variance decreasing in $I_{it}$. The CSV also includes a pre-computed CTR column; throughout this paper we define $CTR_{it}:=C_{it}/I_{it}$ for $I_{it}>0$ and use clicks and impressions for all modelling and evaluation.
@@ 6.1.2 Alignment with empirical magnitudes
The panel is calibrated to resemble plausible ranges reported in industry and marketing analytics contexts: impression volumes are typically in the thousands to tens of thousands per day, and CTR values concentrate in the low single-percent range (approximately 0.5%–3.5%), consistent with broad digital advertising benchmarks (IAB, 2023). The dataset is not intended to reproduce any specific platform, sector, bidding system, targeting policy, or auction environment; rather, it provides a transparent setting for comparing detectors under a clearly specified rule-based evaluation target. The synthetic-only design should be read as proof-of-concept evidence rather than as a substitute for deployment validation.
Importantly, the dataset does not include injected changepoints or labelled fatigue onsets. We therefore define fatigue onset operationally in Section 6.4 and evaluate methods against that rule-based definition. This design supports a proof-of-concept comparison against an explicit operational rule, not a claim of recovery against independently known fatigue events.
@@ 6.2 Comparative Method Analysis
We benchmark our signature-based approach against established methods from both marketing practice and statistical literature. Table 6.2 provides an informal qualitative summary rather than a result established by the experiments in this paper:
@@
Qualitative context for creative fatigue detection methods Method Cost Train Delay Robust. Interpret. Scale Path Signature $O(n)$ No Tunable High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(n\,w_{reg})$ No 2-4 days Low High Good PELT1 $O(n)$ No Offline segmentation Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
@@
Table 6.2 is intended as qualitative context, not as an empirical ranking of all listed methods. The rule-based panel experiment evaluates the signature detector against two practical heuristic baselines because those baselines are common in operational monitoring. The injected-changepoint benchmark additionally includes MOSUM-style adjacent-window mean drops, a GLR-style adjacent-window mean-drop statistic, CUSUM, a hand-crafted window-feature detector, a Gaussian-kernel window-distance detector, and an offline mean-shift reference. A stronger validation would extend this comparison to PELT, Bayesian online changepoint detection, and additional heteroskedastic or exposure-weighted window methods under matched alert-burden constraints.
The signature-based approach offers several practical advantages. Unlike supervised machine learning methods, it requires no labelled training set; however, in the main panel experiment its decision threshold is calibrated retrospectively rather than deployed prospectively on new creatives. Its responsiveness is tunable through the window size and threshold parameter, enabling a practical trade-off between earlier alerts and alert burden. Its fixed-depth complexity is suitable for portfolio screening. These advantages should be interpreted alongside the current experiment’s limitations: the evaluation target is rule-based, the main threshold is retrospective, and the present baseline set is not exhaustive.
@@ 6.3 Experimental Design
    Rule-based evaluation target: Define fatigue onset $\tau$ per creative via the noise-robust, benchmark-relative rule in Section 6.4.
@@
    Method Application: In the panel experiment, apply the signature-based detector and two baseline heuristics (moving-average crossover; rolling regression) to each creative. In the injected benchmark, use the comparator set listed below.
@@
    Injected Benchmark: Run a stochastic simulation with known fatigue onset times and no-event controls, comparing retrospective and past-only signature variants against MOSUM-style mean drops, a GLR-style mean-drop statistic, CUSUM, an offline mean-shift baseline, an exposure-aware signature variant, a hand-crafted window-feature baseline, and a Gaussian-kernel adjacent-window distance baseline.
  6. 6.
    Ablation and stress tests: Vary signature depth and report false-alert rates on no-event nuisance scenarios, including impression shifts, seasonality changes, and delivery gaps without latent CTR fatigue.
  7. 7.
    Operational checks: Report a benchmark-relative lost-clicks proxy under each alerting rule and a runtime smoke benchmark for portfolio screening.
  8. 8.
    Sensitivity Analysis: Vary detector parameters to characterise the trade-off between retrospective timing and alert burden.
For the baseline methods, we use $CTR_{t}:=C_{t}/I_{t}$ and define windows over observed days (i.e., the last $k$ available observations for a creative) to accommodate missingness. Moving-average crossover uses trailing simple moving averages $MA_{7}(t)$ and $MA_{21}(t)$ computed over the last 7 and 21 observed days, respectively. The 21-day moving average and rolling-regression baselines become eligible only once at least 21 observed days are available. An alert triggers on observed day $t$ if $MA_{7}(t)<MA_{21}(t)$ and this inequality holds for three consecutive observed days. Rolling regression fits an unweighted OLS model of $CTR$ on time over the last $w_{reg}=21$ observed days and triggers when the estimated slope is negative and statistically significant at $\alpha=0.05$ for three consecutive observed days. This slope test is a simple heuristic baseline rather than a variance-calibrated test under the binomial noise model; exposure-weighted or heteroskedasticity-robust rolling regressions would be stronger comparators.
The experimental framework supports within-experiment comparison by:
  • using identical data for all methods within each experiment;
  • using consistent pre-processing and smoothing assumptions in the rule-based evaluation target;
  • defining a common timing marker $\hat{\tau}$ for each method as specified in Section 7; and
  • reporting sensitivity analysis over detector parameters to characterise the early-warning versus alert-burden trade-off.
6.4 Operational Definition of Fatigue Onset
Our empirical evaluation requires a reproducible definition of “fatigue onset” that is robust to the substantial day-to-day volatility inherent in digital advertising metrics. Because the dataset does not contain injected changepoints or independently labelled onset dates, we adopt a rule-based evaluation target based on a sustained, economically meaningful deterioration in click-through rate (CTR) relative to a recent performance baseline. This target should not be read as external ground truth; it is an operational proxy used to make the comparison reproducible.